# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ 1 Initialize and dataset $\mathcal{D}_{\text{opt}}=\emptyset$
Output: Curated dataset $\mathcal{D}_{\text{opt}}$, design matrix $V$.
@@ 4.2 Theoretical result
Theorem 4.1.
Instantiating Algorithm 2 with $C>0$ and $\lambda>1$, it holds for any $\delta>0$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
@@
In addition, the number of times the condition of line 4 holds is at most $\frac{d}{\log(1+C)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right)$. Therefore, the size of the batches is on average of order $\tilde{\mathcal{O}}(T/(d\log(T)))$.
Compared to Theorem 3.4, the above regret bound increases only by constant factors. Regarding the number of preference queries, Sekhari et al. (2023) shows that at least $\Omega(T)$ queries are required to achieve $\mathcal{O}(\sqrt{T})$ regret in the worst case. However, optimal design may lead to significantly fewer queries in favorable instances, as demonstrated in our experiments. Importantly, approximate optimal design does not compromise our guarantees, as it ensures a $1/(1-1/e)$ approximation to the optimal.
@@
We present here the key ideas of the proof of Theorem 4.1, the full proof is in Appendix D. The lazy update mechanism used in our algorithm does not degrade the regret bound beyond a constant factor. This follows from the analysis of Abbasi-Yadkori et al. (2011), which shows that the elliptical potentials (related to the determinant of the design matrix) grow by at most a factor of $(1+C)$ compared to the standard design matrix of Algorithm 1. For the experimental design component, we adapt the proof of the standard elliptical lemma (Lattimore and Szepesvári, 2020, Lemma 19.4) to bring out the information gain, measured as $\log\det W_{t}$. This information gain is then related to $\log\det V_{t}$ thanks to standard results from submodular optimization. Specifically, the greedy strategy in Algorithm 3 achieves a $(1-1/e)$ approximation of the optimal information gain. Together, these results allow us to derive a regret bound for this improved algorithm, which matches the original up to constant factors.
@@ 5 Experiments
We validate our theoretical results on regret minimization (Theorems 3.4 and 4.1) on the Isaac-Cartpole-v0 environment from Nvidia Isaac Lab (Mittal et al., 2023). In this task the goal is to balance a pole on a cart by applying left or right forces, preventing the pole from falling. We compare RPO (for $\alpha=1$) with its lazy variants (LRPO without optimal design and LRPO-OD with optimal design). To simulate human preferences, we generate synthetic preferences using the built-in task-specific reward function. Furthermore, we adopt PPO (Schulman et al., 2017) as our RL oracle using 30 PPO steps per iteration of Algorithm 1 and 2. Moreover, to reduce variance in the reward estimate, we sample five trajectories from $\pi_{t}$ and $\pi_{t}^{\prime}$ in each RLHF round.
Figure 1: Comparison of RLHF algorithms in terms of (a) the ground truth reward $V_{\theta^{*}}^{\pi_{t}}$ (estimated from samples) and (b) number of preference queries performed. In particular, for the choice $\alpha=1$, we compare RPO (green, Algorithm 1) with it’s lazy versions LRPO and LRPO-OD (orange & blue). Here, LRPO and LRPO-OD refer to Algorithm 2 without and with optimal design subroutine. The solid lines indicate the median and the shaded areas the 0.2 and 0.8 quantiles, across 10 independent runs. The dashed blue line indicates the mean reward achieved by PPO with the ground truth parameter $\theta^{*}$.
Figure 1 shows that all three algorithms achieve performance competitive with RL using ground truth rewards. However, LRPO-OD reaches the highest performance while requiring significantly fewer preference queries. This highlights that, despite theoretical worst-case lower bounds, the number of preference queries can be considerably reduced in practice by selecting informative queries with optimal design. Additional experimental results and further implementation and evaluation details are provided in Appendix F. Furthermore, our code to run these experiments on Isaac Lab’s manager-based environments is included in the supplementary materials and will be publicly released.
@@ 6 Conclusion
We introduced a simple meta-algorithm for reinforcement learning from human feedback (RLHF) that leverages randomized exploration to achieve both regret and PAC-style guarantees. Our framework applies to general MDPs and utilizes an RL oracle as a subroutine. Notably, to our knowledge, our PAC-RL guarantee is the first for a randomized algorithm in both RLHF and RL settings. Building on this foundation, we developed an improved variant with better scalability and query efficiency, for which we also established regret guarantees. This variant employs lazy updates to enable parallelization of preference oracle calls, significantly reducing the demand on human annotators. Additionally, we apply greedy optimal design to select informative comparisons. Empirically, our approach is competitive with reward-based RL while requiring significantly fewer preference queries. Overall, our contributions advance the state of RLHF by combining strong theoretical guarantees with practical algorithm design, improving efficiency and broadening applicability to real-world scenarios.
Our work opens several directions for future research. First, we provide separate algorithms for regret minimization and PAC guarantees; it remains unclear whether this separation is necessary, and developing a unified algorithm that simultaneously achieves both objectives is an open challenge. Second, we adopt the Bradley-Terry model for preference generation, which may not fully capture the complexity of real-world human feedback. Extending the framework to richer preference models is an important direction. Third, our approach relies on an RL oracle at every RLHF step, which may be computationally demanding. While using a reward-free algorithm as an RL oracle is theoretically efficient, practical RL implementations are typically based on policy optimization, which is not a reward-free algorithm, and often entails high sample complexity. Thus, it remains open whether we could require fewer RL oracle calls or whether reward-free oracles can be successfully implemented. Finally, our experimental evaluation is limited to simple robotic control tasks with synthetic feedback. Assessing performance on more complex tasks and with real human or LLM-generated feedback would offer a stronger test of the method’s practical applicability.

### v2

@@ 1 Initialize dataset $\mathcal{D}_{\text{opt}}=\emptyset$
Output: $(\mathcal{D}_{\mathrm{opt}},V)$ if $\det(V)\geq\det(W)$, else $(\mathcal{D},W)$
@@ 4.2 Theoretical result
{restatable}
theoremlazy Instantiating Algorithm 3 with $C>0$ and $\lambda\geq 4H_{\gamma}^{2}L^{2}$, it holds for any $\delta\in(0,1)$ and $T\in\mathbb{N}$, with probability at least $1-\delta$ that
@@
In addition, the number of times the condition of line 4 holds is at most $\frac{d}{\log(1+C)}\log\left({1+\frac{T(2LH_{\gamma})^{2}}{d\lambda}}\right)$. Therefore, the size of the batches is on average of order $\tilde{\mathcal{O}}(T/(d\log(T)))$.
Compared to Theorem 3.2.1, the above regret bound increases only by constant factors. Regarding the number of preference queries, sekhari2023contextual shows that at least $\Omega(T)$ queries are required to achieve $\mathcal{O}(\sqrt{T})$ regret in the worst case. However, optimal design may lead to significantly fewer queries in favorable instances, as demonstrated in our experiments, while preserving the worst-case guarantees.
@@
We briefly outline the main idea of the proof of Theorem 4.2; the full argument is deferred to Appendix C. The central observation is that similarly as for the standard lazy update analysis (abbasi2011improved), the regret only increases by a constant factor $(1+C)$ given that the optimal design subroutine ensures that $\det V_{t_{\operatorname{stop}}}\geq\det W_{t_{\operatorname{stop}}}$ and $|\mathcal{D}_{\text{opt}}|\leq|\mathcal{D}|$.
@@ 5 Experiments
We first validate our theoretical results on regret minimization in a tabular gridworld environment, where our RL oracle assumption provably holds, and then compare Algorithm 1 and 3 on more challenging continuous control tasks. The validation of Algorithm 2 is deferred to Appendix F.
5.1 Tabular environment
Figure 1: We compare the regret of RPO-Regret (orange, Algorithm 1) against a baseline with entropy exploration (blue). The solid lines indicate the median and the shaded areas the 0.2 and 0.8 quantiles, across 20 independent runs. The regret is computed with respect to the ground truth reward parameter $\theta^{*}$.
We consider a $6\times 6$ gridworld environment with deterministic transitions and $4$ actions (up, down, left, right). The agent starts in the center and receives reward $0.5$ if it reaches one of two boundary states. The reward features are one-hot features of six boundary states – including the two goal states (see Figure 3 for an illustration of the environment). We compare RPO-Regret (Algorithm 1) against an RLHF baseline that explores purely through entropy regularization, using synthetic preferences generated from the ground truth reward parameter $\theta^{*}$. We compute optimal policies using soft value iteration (haarnoja2017reinforcement), and to reduce variance in the reward estimate, we sample $100$ trajectories from $\pi_{t}$ and $\pi_{t}^{\prime}$ in each RLHF round.
As shown in Figure 1, RPO-Regret attains considerably lower regret with less variance across runs than the baseline. These results underscore that entropy exploration, does not necessarily guarantee low regret in MDPs.
5.2 Continuous control
Figure 2: Comparison of regret minimization algorithms in terms of (a) the cumulative regret (estimated from samples against an RL policy trained with $\theta^{*}$) and (b) number of preference queries performed. In particular, we compare RPO-Regret (orange, Algorithm 1) with its lazy versions LRPO-Regret and LRPO-OD-Regret (blue & green). Here, LRPO-Regret and LRPO-OD-Regret refer to Algorithm 3 without and with optimal design subroutine. The solid lines indicate the median and the shaded areas the 0.2 and 0.8 quantiles, across 20 independent runs.
We validate our theoretical results on regret minimization (Theorems 3.2.1 and 4.2) on the Isaac-Cartpole-v0 environment from Nvidia Isaac Lab (mittal2023orbit). In this task the goal is to balance a pole on a cart by applying left or right forces, preventing the pole from falling. We compare RPO-Regret with its lazy variants (LRPO-Regret without optimal design and LRPO-OD-Regret with optimal design). To simulate human preferences, we generate synthetic preferences using the built-in task-specific reward function, which is a linear combination of $5$ reward terms (see Appendix F). Furthermore, we adopt PPO (schulman2017proximal) as our RL oracle using 30 PPO steps per iteration of Algorithm 1 and 3. Again, we sample $100$ trajectories from $\pi_{t}$ and $\pi_{t}^{\prime}$ in each RLHF round.
Figure 2 shows that for all three algorithms the regret slope flattens after a few RLHF rounds, demonstrating sublinear regret and performance competitive with RL using ground truth rewards. Although LRPO-OD-Regret suffers slightly higher regret initially, it quickly reaches the same performance as RPO-Regret and LRPO-Regret, while reducing the number of preference queries considerably. This highlights that, despite theoretical worst-case lower bounds, the number of preference queries can be considerably reduced in practice by selecting informative queries with optimal design. Additional experimental results and further implementation and evaluation details are provided in Appendix F.
@@ 6 Conclusion
We presented two simple meta-algorithms for reinforcement learning from human feedback (RLHF) that combine an RL oracle and randomized exploration, achieving complementary guarantees: regret bounds for Algorithm 1 and PAC-style guarantees for Algorithm 2. Algorithm 2, to our knowledge, is the first tractable method to perform preference-free exploration with trajectory-level preferences. Building on this framework, we introduced Algorithm 3, a practical regret-minimization algorithm that combines lazy updates to enable parallelization with optimal design to reduce query complexity, while maintaining regret guarantees. Empirically, our approach is competitive with reward-based RL while requiring significantly fewer preference queries. Overall, our contributions advance the state of RLHF by combining strong theoretical guarantees with practical algorithm design, improving efficiency and broadening applicability to real-world scenarios.
Our work opens several directions for future research. First, we adopt the Bradley-Terry model for preference generation, which may not fully capture the complexity of real-world human feedback. Extending the framework to richer preference models is an important direction. Second, our approach relies on an RL oracle at every RLHF step, which may be computationally demanding. While using a reward-free algorithm as an RL oracle is theoretically efficient, practical RL implementations are typically based on policy optimization, which is not a reward-free algorithm, and often entails high sample complexity. Thus, it remains open whether we could require fewer RL oracle calls or whether reward-free oracles can be successfully implemented. Finally, our experimental evaluation is limited to tabular settings and simple robotic control tasks with synthetic feedback. Assessing performance on more complex tasks and with real human or LLM-generated feedback would offer a stronger test of the method’s practical applicability.
Acknowledgments
Andreas Schlaginhaufen is funded by a PhD fellowship from the Swiss Data Science Center.